OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope (–7, 9); m = 4/5

OpenStudy (anonymous):

The equation in general point-slope form is:\[y - y _{1} = m(x - x _{1})\]where "m" is your slope and\[(x _{1}, y _{1})\]is the point you are given. So now, just substitute.

OpenStudy (anonymous):

If my options are y + 7 = (x – 9) y – 7 = (x + 9) y – 9 = (x + 7) y + 9 = (x – 7) Would it be c ?

OpenStudy (anonymous):

What happened to the slope of 4/5? That doesn't seem to be appearing in any of the selections. Your selection "c" is the closest but it needs a slope in front of the parentheses.

OpenStudy (anonymous):

I know, this question is rather confusing. I will go with C, I suppose. Thank you for your help!

OpenStudy (anonymous):

You're welcome! Good luck in all of your studies and thx for the recognition!